OpenStudy (sugarmochi):

P is directly proportional to Q when P is 22, Q is 4. Find the equation connecting P and Q Hence find : P when Q=7

OpenStudy (mathmate):

Hint: start with the relation P=kQ where k is a constant to be found using the fact that P=22 when Q=k.

OpenStudy (eliesaab):

You mean when P=22 and Q=4

OpenStudy (eliesaab):

22 = 4 k How much is k?

OpenStudy (eliesaab):

Are you with me?

OpenStudy (mathmate):

@SugarMochi @eliesaab is helping you, please respond. @eliesaab yes, I meant Q=4, that makes 22=4k as you had it.

OpenStudy (sugarmochi):

Im-- yeah im following somehow

OpenStudy (eliesaab):

Can you solve 22=4 k for k?

OpenStudy (eliesaab):

\[ k=\frac{22}4=\frac {11}2\]

OpenStudy (eliesaab):

\[ P= \frac {11}2 Q =\frac {11}2 7=\frac {77}2=38.5 \]

OpenStudy (mathmale):

@eliesaab: If you're going to ask someone, "can you solve so and so for k," give the other person time to do it. Why were you doing someone else's assignment?

OpenStudy (mathmale):

Your job is to enable the other person to find his or her own answer, that is, to do his or her own work.